OpenStudy (anonymous):

one quick question, medal awarded!!!!!!! Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 to the 3 over 4 power, all over 2 to the 1 over 2 power the eighth root of 2 to the third power the square root of 2 to the 3 over 4 power the fourth root of 2 the square root of 2

OpenStudy (solomonzelman):

\[\sqrt[c]{a^b}=a^{b/c}\]

OpenStudy (mathstudent55):

This is how you solve the first one. \(\dfrac{a^m}{a^n} = a^{m - n} \)